# The Future of Large Language Model Pre-training is Federated

### v1

@@ 4 Evaluation
Table 1: Our scalability experiments’ model sizes and afferent hyperparameter settings. We use common parameters for the AdamW [75] optimizer for the local training. We reset the local optimizer states every round as it helps convergence after aggregation. The centralized model uses the same local hyperparameters.
@@
For all our experiments, we use a version of C4 [76] randomly split into $8$ equally-sized clients trained across a variable number of rounds depending on the model size. All experiments use $500$ local steps per round executed by all the clients in the federation. We distinguish between parallel steps and sequential steps since the model parameters are averaged after each round, and each of the $500$ local steps done during that round is done in parallel, as in standard data-parallel training [34]. Our setting respects the conditions of standard cross-silo [50] FL in which the orchestrator samples the entire batch of equally capable clients in every round. Our experiments used diverse hardware resources whose specifications allow for a reasonably fast execution. Our models have been trained on a combination of heterogeneous servers equipped with NVIDIA A40, A100, and H100 GPUs. Due to FL’s lax synchronization requirements, these heterogeneous hardware accelerators could collaborate despite being located in different countries.
@@
Our experiments focused on scaling the model size to determine if federated optimization can train models approximating the size of those commonly used today [29, 30]. As shown in Table 1, we trained models ranging in size from $75$ million parameters to $1.3$ billion.
We observed that the stability of the training procedure improves directly with the model size, with federated optimization reaching a better consensus across client models. As shown in Fig. 2, while the centralized model’s final performance is slightly greater for small model sizes, it becomes identical for the $1.3$B model. Furthermore, client training perplexities showcase a shift from a transient oscillatory phase to a convergence phase where federated optimization acts as a regularizer, improving performance. Thus, while aggregation increases local perplexity in the early stages, it decreases it at convergence. This transition is much quicker for larger models, with the $1.3$B spending only $4$ rounds in the pre-convergence phase, while the $75$M requires over $20$ rounds to reach a consensus across client models. Similarly, while the $75$M model continues to experience occasional turbulences in convergence after round $20$, the larger models, such as the $350$M, obtain full client model alignment.
Figure 2: Perplexity comparison between the server model evaluated on the centralized validation set, the train and test perplexities for a centralized baseline, and the training perplexities of clients for our $75$M (a), $125$M (b), $350$M (c), and $1.3$B experiments. Crucially, the stability of federated training increases with model size. For example, the centralized model outperforms the $75$M federated while performing near-identically for the $1.3$B case. While federated aggregation initially causes large spikes in client perplexity, these subside as the clients reach a consensus on the model parameters, which happens much quicker for larger models. Following this transitory phase, aggregation applies a regularizing effect on the model performance, allowing a better model to be trained than would be possible for a single client. The server validation perplexity is a soft upper bound for the spikes, with the gap between train and validation perplexities decreasing over time.
@@
We further investigate the training dynamics of the models of size $75$ and $350$ million trainable parameters. In particular, we are interested in understanding how the client and server models interplay as the FL converges. As Fig. 3 shows, the server model initially “pulls back” the clients’s model norms through the aggregation. After a few rounds, the federated aggregation starts incrementing the norm of the averaged clients’ models until the global and local models converge to the same norm. Figure 3 showcases how the client and server optimizations interplay to agree on an initial set of global model parameters and then converge to an optimum.
Figure 3: The $l_{2}$ norms of the global model, client models, and the average of client models for our $75$M (a) and $350$M (b) experiments. While in the transitory early stage, the server model grows slower than the aggregate of client models, it is consistently larger in later stages. This reflects both clients reaching a consensus and the ability of the momentum mechanism to stabilize the optimization trajectory.
@@
The cause of client model convergence can be observed in Fig. 4, which showcases the relationship between the norms of the per-round pseudo-gradient (i.e, average client update) and the local client gradients applied at every SGD step. While the pseudo-gradient starts much larger than the applied local gradients, it reaches a similar or much smaller size as the client models converge.
Figure 4: The $l_{2}$ norms, for our $75$M (a) and $350$M (b) experiments, of the FedAvg Pseudo-Gradient (average of client deltas relative to the server model of the previous round), the client model gradients computed on a per-step basis, and the client gradients applied to the model when considering learning rate, weight decay, and clipping. The decay in the FedAvg pseudo-gradient is much faster than in the local gradients. We postulate that averaging cancels conflicting and potentially noisy optimization directions.
@@ 6 Conclusions
We successfully demonstrate the potential of LLM federated generative pre-training by obtaining the first federated billion-scale model fully trained in a heterogeneous FL setting. The complete system presented in this work allows for collaborative, reproducible, and scalable pre-training of LLMs, releasing the computational and data resources spread across the planet. Our system is built on an open-source framework [23] and will soon be made publicly available. Furthermore, we fully disclose our training recipe to enable future development and collaborations. Our work advances LLM pre-training by democratizing it through federated technologies, allowing for data-rich actors to pool together resources despite their individual compute and connectivity capabilities. Given current trends, we believe that the future of LLM pre-training lies in the data sources that FL can bring together.

### v2

@@ 6.3 Experimental Design
6.4 Training Setup
@@
All experiments use $500$ local steps per round, executed by all the clients in the federation. We distinguish between parallel steps and sequential steps since the model parameters are averaged after each round, and each of the $500$ local steps in a round is done in parallel, as in standard data-parallel training [34]. Our setting respects the conditions of standard cross-silo [50] FL in which the orchestrator samples the entire batch of equally capable clients in every round. Our experiments used diverse hardware resources whose specifications allow for reasonably fast execution; as an example, we used a combination of heterogeneous servers equipped with NVIDIA A40, A100, and H100 GPUs. Due to FL’s lax synchronization requirements, these heterogeneous hardware accelerators were able to collaborate despite being located in different countries.
As shown in Table 1, we trained models ranging in size from $75$ million parameters to $1.3$ billion for the causal language modeling task. We used the tokenizer presented in [101] with a vocabulary size of 50 3685036850\,36850 368. The local optimizer the clients use in our experiments is AdamW [102], while the server optimizer is FedMom [47]. The hyperparameters we used are reported in Table 2.
Table 2: Hyperparameters used in our experiments. The federated learning rate $\boldsymbol{\eta_{s}}$ and momentum $\boldsymbol{\mu_{s}}$ [47] are applied by Photon Aggregator. The FL is performed across the number of rounds reported. $\mathbf{S_{C}}$ are the parameters of the learning rate scheduler synchronized across sequential steps. $\alpha$ is the factor to be applied to the maximum learning rate $\eta_{max}$ to obtain the minimum learning rate for the cosine scheduler, i.e., $\eta_{min}=\alpha\times\eta_{max}$. $T$ is the duration, in steps, of the cosine scheduler. We also report the batch size used in the local training by the Photon LLM Nodes.
Model Size #Rounds $\boldsymbol{\eta_{s}}$ $\boldsymbol{\mu_{s}}$ $\boldsymbol{\alpha}$ $\boldsymbol{\eta_{max}}$ $\boldsymbol{T}$ Batch Size
$\mathbf{75}$M $40$ $0.2$ $0.9$ $10^{-6}$ $4\times 10^{-4}$ 88 0008800088\,00088 000 $256$
$\mathbf{125}$M $25$ $0.2$ $0.9$ $10^{-5}$ $3\times 10^{-4}$ 15 0001500015\,00015 000 $256$
$\mathbf{350}$M $40$ $0.2$ $0.9$ $10^{-1}$ $3\times 10^{-4}$ 13 4001340013\,40013 400 $256$
$\mathbf{1.3}$B $16$ $0.2$ $0.9$ $10^{-1}$ $2\times 10^{-4}$ 24 8002480024\,80024 800 $512$
7 Evaluation
We have experimentally validated the efficacy of our federated training recipe at a series of increasing model scales while investigating impactful choices for both the federated and local optimization. Thus, the results presented in this section indicate the viability of federated optimization for LLM pre-training and validate its ability to reach beyond the data and hardware constraints of centralized approaches.
Figure 3: Comparison between the perplexity of the federated global model evaluated on the centralized validation set, the train perplexities of federated clients (averaged together), and the train and test perplexities for a centralized experiment. These metrics are reported for our $75$M (a), $125$M (b), $350$M (c), and $1.3$B (d) experiments. Crucially, the stability of federated training increases with model size. For example, the centralized model outperforms the $75$M federated model while performing near-identically for the $1.3$B models. While federated aggregation initially causes large spikes in client perplexity, these subside as the clients reach a consensus on the model parameters, which happens much quicker for larger models. Following this transitory phase, aggregation applies a regularizing effect on the model performance, allowing a better model to be trained than would be possible for a single client. The server validation perplexity is a soft upper bound for the spikes, with the gap between train and validation perplexities decreasing over time.
Federated Optimization is Effective and Stable across Scales:
Our initial experiments using the IID partition of the C4 dataset reveal that federated optimization can reach a performance comparable to centralized training. As can be seen in Fig. 3, federated models perform comparatively well to centrally trained ones or even better at sufficiently large model sizes. Furthermore, while their convergence curves may initially look more unstable, the federated aggregation procedure quickly resolves this once the model settles into an appropriate compromise for the clients. Most importantly, FL effectively trains the $\mathbf{1.3}$B parameters model. The gap between centralized and federated performance becomes increasingly small as the model size increases, ranging between $10$ and $2$ points on the perplexity axis.
As such, we believe that federated optimization can effectively replace standard data-parallel techniques in data center settings when nodes are poorly connected, which Photon handles automatically. Furthermore, we will now argue that it is suitable for in-the-wild federated systems where data may also come from statistically heterogeneous distributions.
Federated Optimization is Robust to Heterogeneity:
As previously discussed, federated statistical heterogeneity problem is inevitable for real federated learning scenarios as clients use naturally generated private data to train their models. Thus, we investigate how much heterogeneity impacts performance using our federated partition of The Pile composed of Wikipedia (en), ArXiv, Project Gutenberg (PG-19), HackerNews, PubMed Central, FreeLaw, PhilPapers, and StackExchange. We conduct these experiments on our smaller $75$M and $125$M models. We chose them because they are under-parameterized compared to our $1.3$B model, making reconciling heterogeneous data particularly challenging. In Table 2, we report the hyperparameters we use for our experiments.
We observe a high degree of robustness towards this naturally heterogeneous partition. As shown in Fig. 4, while clients show much more significant variance in their loss, the server validation loss converges with a similar trend to the IID C4 case. Compared to the previous set of experiments, the absolute performance improves from 50 to 35 for the $75$M model and from 50 to 30 for the $125$M model mainly because of the different datasets used. We believe that the highly effective convergence is due to the regularization properties of FL, precisely because of FL’s meta-learning properties [103, 104, 105].
Furthermore, as can be seen in Fig. 5, models trained in a federated fashion are not subject to the same tendency of rapidly increasing activations as centralized models are. Corroborating this against the perplexity in Fig. 4 shows that while the centralized model may diverge beyond repair, the federated model can quickly recover from temporary decreases in server validation accuracy. We argue that this is not a mere artifact of the model being less trained since Fig. 5 show that activations of the centralized model start higher than those of the federated one and remain so throughout the entire duration of the experiments. We attribute this to the noise injected by the aggregation procedure, given that we observe periodic reductions in activation norms at round boundaries. This confirms the robustness properties of federated aggregation [18].
Figure 4: Perplexity comparison between the global model evaluated on the centralized validation set, the train and test perplexities for a centralized baseline, and the training perplexities of clients (averaged together) for our naturally heterogeneous partition of The Pile using either a $75$M (a) or $125$M (b) model size. Unlike the homogeneous partition, the natural heterogeneity of the underlying datasets makes an initial consensus harder to reach for the federated model, as can be observed from the very high initial clients and server perplexities. However, like the IID partition shown in Fig. 3, once clients reach consensus, performance becomes comparable to a centralized baseline.
Figure 5: The $l_{2}$ norms of the output activations of our $75$M (a) and $125$M  (b) models trained on a naturally heterogeneous partition of The Pile. The norm of the activations is a well-known indicator of future model divergence [94]. As can be observed, the activations of the centralized model outpace those of the federated clients right from the start and experience a massive increase towards the end of training. The aggregation procedure keeps the federated clients in check by reducing the norm of the activations round to round.
@@ 7 Evaluation
Our experiments focused on scaling the model size to determine if federated optimization can train models approximating the size of those commonly used today [29, 30].
We observed that the stability of the training procedure improves directly with the model size, with federated optimization reaching a better consensus across client models. As shown in Fig. 3, while the centralized model’s final performance is slightly greater for small model sizes, it becomes almost identical for the $1.3$B model. Furthermore, client training perplexities showcase a shift from a transient oscillatory phase to a convergence phase where federated optimization acts as a regularizer, improving performance. Thus, while aggregation increases local perplexity in the early stages, it decreases it at convergence. This transition is much quicker for larger models, with the $1.3$B model spending only $4$ rounds in the pre-convergence phase, while the $75$M model requires over $20$ rounds to reach a consensus across client models. Similarly, while the $75$M model continues to experience occasional turbulences in convergence after round $20$, the larger models, such as the $350$M model, obtain full client model alignment.
Federated Optimization is Robust to Partial Participation:
Our experiments have shown that meaningfully-sized LLMs can be trained in a federated fashion when using full client participation. However, one particularly effective way of reducing overall training costs in terms of parallel computing while still having the regularization benefits of training on more data is to subsample a smaller portion of the entire federated population. The experiments in Fig. 6 show how subsampling only $6.25\%$ of clients from our $64$-client partition results in the same performance as our previous full-participation experiments, despite training only $4$ clients per round instead of $8$. Such robustness to partial participation opens the door to highly efficient LLM training schedules in which the scalable addition of data, as needed, improves convergence whilst minimizing computing resource utilization. Furthermore, it enables applications where multiple federated workloads may be executed simultaneously by sampling different clients from the population.
Figure 6: Perplexity comparison, $75$M (a) and $125$M (b), between the server model evaluated on the centralized validation set, the train and test perplexities for a centralized baseline, and the training perplexities of clients for our $64$-client partition of C4 with $4$ client participating per round for a $6.25\%$ participation rate. Despite using half of the parallel compute of the full participation baselines, partial participation converges to the same performance at the expense of more turbulences early in the training process. Thus, subsampling cohorts can substantially reduce hardware utilization for a sufficiently large cross-silo population, allowing for parallel training of several models.
@@
We further investigate the training dynamics of the models of size $75$M and $350$M parameters. In particular, we are interested in understanding how the client and server models interplay as the FL converges. As Fig. 7 shows, the server model initially “pulls back” the clients’ model norms through the aggregation. After a few rounds, the federated aggregation starts incrementing the norm of the averaged clients’ models until the global and local models converge to the same norm. Figure 7 showcases how the client and server optimizations interplay to agree on an initial set of global model parameters and then converge to an optimum. The same trends are present for our heterogeneous partition of The Pile (Fig. 10) and for our partial-participation experiments (Fig. 11.
Figure 7: The $l_{2}$ norms of the global model, client models, and the average of client models for our $75$M (a) and $350$M (b) experiments. While in the transitory early stage, the server model grows slower than the aggregate of client models, it is consistently larger in later stages. These figures reflect both clients reaching a consensus and the ability of the momentum mechanism to stabilize the optimization trajectory.
@@
We observe the cause of client model convergence in Fig. 8, which showcases the relationship between the norms of the per-round pseudo-gradient (i.e., average client update) and the local client gradients applied at every SGD step. While the pseudo-gradient starts much larger than the applied local gradients, it reaches a similar or much smaller size as the client models converge. The same trends hold for our results on heterogeneous data, shown in Fig. 12, and for our partial-participation experiments, shown in Fig. 13.
Figure 8: The $l_{2}$ norms, for our $75$M (a) and $350$M (b) experiments, of the FedAvg Pseudo-Gradient (average of client deltas relative to the server model of the previous round), the client model gradients computed on a per-step basis, and the client gradients applied to the model when considering learning rate, weight decay, and clipping. The decay in the pseudo-gradient is faster than the decay of the scheduler for the $75$M model, being data-driven, and comparable to that of the step-gradient for the $125$M model.
@@ 9 Conclusions
We successfully demonstrate the potential of LLM federated generative pre-training by obtaining the first federated billion-scale model fully trained in a heterogeneous FL setting and showing how federated training is robust to both data heterogeneity and partial participation. The complete system presented in this work allows for collaborative, reproducible, and scalable pre-training of LLMs, releasing the computational and data resources spread across the planet. Our system is built on an open-source framework [23] and will soon be made publicly available. Furthermore, we fully disclose our training recipe to enable future development and collaborations. Our work advances LLM pre-training by democratizing it through federated technologies. This allows data-rich actors to pool resources together, whether extensive or limited. Given current trends, we believe that the future of LLM pre-training lies in the data sources that FL can bring together.
Acknowledgments and Disclosure of Funding
All costs for the computation used for this work was funded by Flower Labs, and the research conducted by a team of researchers from Flower Labs and The University of Cambridge. Support for university based researchers came from a variety of sources, but in particular the following funding organizations are are acknowledged: the European Research Council (REDIAL), the Royal Academy of Engineering (DANTE) and the Ministry of Education of Romania through the Credit and Scholarship Agency.
Appendix A Additional Results
Figure 9: The $l_{2}$ norm of the global model versus the server-side Nesterov momentum, which seeks to track an exponential moving average of the model using $\beta=0.7$, for our 75M (a), 125M (b), 350M (c), and 1.3B (d) experiments.
Figure 10: The $l_{2}$ norms of the global model, client models, and the average of client models for our $75$M (a) and $125$M (b) experiments on our naturally heterogeneous partition of The Pile. Despite the heterogeneous data, the momentum mechanism of the server and the decaying learning rate allow it to enforce a consensus effectively.
Figure 11: The $l_{2}$ norms of the global model, client models, and the average of client models for our $75$M (a) and $125$M (b) experiments on our partial participation experiments using $4$ clients per round out of $64$ total C4 clients, i.e., the $6\%$ of the federated population. Despite the FL using only a small sample of clients every round, the global model follows the exact same trend as in the full participation scenario.
Figure 12: The $l_{2}$ norms, for our $75$M (a) and $125$M (b) experiments on The Pile, of the FedAvg Pseudo-Gradient (average of client deltas relative to the server model of the previous round), the client model gradients computed on a per-step basis, and the client gradients applied to the model when considering learning rate, weight decay, and clipping. Unlike the IID case, the FedAvg Pseudo-Gradient decays much faster than local client gradients, indicating that its decrease results from the model adapting to data heterogeneity rather than being a consequence of learning rate decay.
Figure 13: The $l_{2}$ norms, for our $75$M (a) and $125$M (b) partial participation experiments on C4, of the FedAvg Pseudo-Gradient (average of client deltas relative to the server model of the previous round), the client model gradients computed on a per-step basis, and the client gradients applied to the model when considering learning rate, weight decay, and clipping. The relationship between the FedAvg Pseudo-Gradient and the local steps is the same as in the full-participation case despite a very small sample of clients being considered every round.